# CLIP Meets Diffusion: A Synergistic Approach to Anomaly Detection

### v1

@@ Appendix G Further results in the one-for-all paradigm
| CLIPFUSION (Ours) | 96.4 | 98.2 | 88.3 | 90.0 | |
Appendix H Detailed quantitative results
Category k=0 k=1 k=2 k=4 WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION Bottle 99.20.0 100.00.0 97.40.7 99.40.4 98.20.9 100.00.0 98.51.0 99.20.3 99.30.3 100.00.0 98.80.2 99.20.3 99.30.4 100.00.0 Cable 86.50.0 87.30.0 57.74.6 88.84.2 88.91.9 95.00.9 62.35.9 91.02.7 88.40.7 96.30.4 70.06.1 91.02.7 90.90.9 96.70.4 Capsule 72.90.0 88.10.0 57.77.3 67.82.9 72.36.8 83.66.9 64.33.0 72.87.0 77.38.8 89.02.7 65.22.5 72.87.0 82.38.9 91.21.6 Carpet 100.00.0 99.90.0 96.61.0 95.30.8 99.80.3 100.00.0 97.80.5 96.60.5 99.80.3 100.00.1 97.90.4 96.60.5 100.00.0 100.00.0 Grid 98.80.0 99.90.0 54.26.7 63.610.3 99.50.3 99.40.4 67.24.2 67.78.3 99.40.2 99.60.4 68.13.8 67.78.3 99.60.1 99.50.4 Hazelnut 93.90.0 96.90.0 88.32.6 88.32.7 97.51.4 99.30.4 90.80.8 93.23.8 98.30.7 99.30.3 91.91.2 93.23.8 98.40.4 99.30.5 Leather 100.00.0 99.90.0 97.50.7 97.30.7 99.90.0 99.90.0 97.50.9 97.90.7 99.90.0 100.00.0 98.50.2 97.90.7 100.00.0 100.00.0 Metal nut 97.10.0 98.30.0 53.03.8 73.42.9 98.70.8 99.80.3 54.83.8 77.78.5 99.40.2 99.90.2 60.75.2 77.78.5 99.50.2 100.00.0 Pill 79.10.0 77.60.0 61.33.8 81.92.8 91.22.1 96.10.4 59.16.4 82.92.9 92.30.7 96.20.2 54.92.7 82.92.9 92.81.0 96.40.4 Screw 83.30.0 77.30.0 55.02.5 44.44.6 86.40.9 72.58.5 54.04.4 49.03.8 86.02.1 73.55.1 50.04.1 49.03.8 87.91.2 82.04.5 Tile 100.00.0 99.10.0 92.22.2 99.00.9 99.90.0 99.90.2 93.31.1 98.51.0 99.90.2 99.90.1 93.10.6 98.51.0 99.90.1 100.00.0 Toothbrush 87.50.0 96.90.0 82.51.2 83.33.8 92.24.9 98.91.6 87.64.2 85.93.5 97.51.6 99.10.8 89.22.5 85.93.5 96.72.6 98.50.8 Transistor 88.00.0 86.70.0 73.36.0 78.16.9 83.43.8 90.91.4 72.86.3 90.04.3 85.31.7 91.10.8 82.46.5 90.04.3 85.72.5 92.91.0 Wood 99.40.0 96.70.0 96.11.2 97.80.3 99.90.1 98.90.3 96.90.5 98.30.6 99.90.1 98.60.2 97.00.2 98.30.6 99.80.3 98.80.1 Zipper 91.50.0 94.00.0 85.82.7 92.30.5 88.85.9 96.50.7 86.32.6 94.02.1 94.01.4 95.32.2 88.32.0 94.02.1 94.50.5 96.20.6 Mean 91.80.0 93.20.0 76.63.1 83.43.0 93.12.0 95.40.8 78.93.1 86.33.3 94.41.3 95.90.3 80.42.5 88.82.6 95.21.3 96.80.2
Category k=0 k=1 k=2 k=4 WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION Bottle 99.80.0 100.00.0 99.20.2 99.80.1 99.40.3 100.00.0 99.60.3 99.80.1 99.80.1 100.00.0 99.70.0 99.80.1 99.80.1 100.00.0 Cable 91.20.0 92.70.0 64.93.8 93.82.2 93.21.1 97.10.5 69.66.6 95.11.3 92.90.6 97.80.3 76.15.6 97.10.7 94.40.3 98.10.2 Capsule 91.50.0 97.20.0 86.92.2 89.42.0 91.62.7 95.92.2 88.40.8 91.02.9 93.33.6 97.40.8 87.80.8 94.91.1 95.13.3 98.00.4 Carpet 100.00.0 100.00.0 99.00.2 98.70.2 99.90.1 100.00.0 99.40.1 99.00.1 99.90.1 100.00.0 99.40.1 98.80.2 100.00.0 100.00.0 Grid 99.60.0 100.00.0 75.03.3 81.14.9 99.90.1 99.80.1 82.52.3 84.14.0 99.80.1 99.90.1 83.01.8 86.44.0 99.90.0 99.80.1 Hazelnut 96.90.0 98.30.0 93.31.7 92.92.2 98.60.7 99.60.2 94.10.5 96.02.0 99.10.4 99.60.1 94.80.6 97.01.2 99.10.2 99.60.3 Leather 100.00.0 100.00.0 99.20.2 99.10.2 100.00.0 100.00.0 99.20.3 99.30.2 100.00.0 100.00.0 99.60.1 99.60.1 100.00.0 100.00.0 Metal nut 99.30.0 99.60.0 82.02.7 91.01.1 99.70.2 99.90.1 82.21.4 92.34.0 99.90.0 100.00.0 85.51.7 97.02.6 99.90.1 100.00.0 Pill 95.70.0 94.80.0 88.31.3 96.50.6 98.30.5 99.30.1 87.92.6 96.60.7 98.60.1 99.30.0 87.01.2 96.90.4 98.60.2 99.40.1 Screw 93.10.0 89.30.0 78.11.0 71.42.3 94.20.6 87.54.7 77.31.3 72.93.4 94.11.5 87.43.2 75.72.8 71.81.9 94.90.8 93.02.3 Tile 100.00.0 99.60.0 97.20.7 99.60.3 100.00.0 100.00.1 97.60.4 99.40.4 100.00.1 100.00.0 97.60.2 99.60.1 100.00.0 100.00.0 Toothbrush 95.60.0 98.80.0 93.70.5 93.51.4 96.72.0 99.60.5 95.21.6 94.11.4 99.00.6 99.70.3 95.80.7 94.80.7 98.71.1 99.50.2 Transistor 87.10.0 84.70.0 66.27.5 77.75.5 79.04.0 89.81.5 69.06.5 89.33.9 80.72.3 89.90.7 77.68.4 84.59.0 80.73.2 92.31.3 Wood 99.80.0 99.00.0 98.80.3 99.30.1 100.00.0 99.70.1 99.00.1 99.50.2 100.00.0 99.60.1 99.10.0 99.50.2 99.90.1 99.70.0 Zipper 97.50.0 97.90.0 95.50.9 97.20.3 96.81.8 98.90.3 95.41.0 97.81.0 98.30.4 98.31.0 96.20.8 99.10.7 98.50.2 98.70.2 Mean 96.50.0 96.80.0 88.11.7 92.21.5 96.50.9 97.80.4 89.31.7 93.81.7 97.00.7 97.90.2 90.51.6 94.51.5 97.30.6 98.50.1
Category k=0 k=1 k=2 k=4 WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION PaDiM PatchCore WinCLIP CLIPFUSION Bottle 89.50.0 94.10.0 96.10.5 97.90.1 97.50.2 97.30.2 96.90.1 98.10.0 97.70.1 97.50.0 97.10.1 98.20.0 97.80.0 97.70.1 Cable 77.00.0 73.20.0 88.41.2 95.50.8 93.80.6 91.30.2 90.00.8 96.40.3 94.30.4 92.20.3 92.10.4 97.50.3 94.90.1 92.90.4 Capsule 86.90.0 94.60.0 94.50.6 95.60.4 94.60.8 97.90.4 95.20.5 96.50.4 96.40.3 98.40.2 96.20.4 96.80.6 96.20.5 98.50.1 Carpet 95.40.0 99.30.0 97.80.2 98.40.1 99.40.0 99.40.1 98.20.0 98.50.1 99.30.0 99.50.0 98.40.0 98.60.1 99.30.0 99.50.0 Grid 82.20.0 98.90.0 70.22.8 58.84.9 96.81.0 98.80.2 70.82.0 62.63.2 97.70.8 98.90.4 77.01.8 69.41.3 98.00.2 99.20.1 Hazelnut 94.30.0 96.70.0 95.40.7 95.80.6 98.50.2 98.20.1 96.80.3 96.30.6 98.70.1 98.20.1 97.20.2 97.60.1 98.80.0 98.70.1 Leather 96.70.0 99.20.0 98.50.1 98.80.2 99.30.0 99.30.0 98.70.1 99.00.1 99.30.0 99.30.0 98.80.0 99.10.0 99.30.0 99.40.0 Metal nut 61.00.0 74.20.0 74.61.1 89.31.4 90.00.6 86.21.2 80.32.1 94.61.4 91.40.4 89.70.9 82.73.9 95.91.8 92.90.4 91.30.8 Pill 80.00.0 83.00.0 84.81.0 93.11.1 96.40.3 95.50.2 87.30.7 94.20.3 97.00.2 95.60.2 88.90.5 94.80.4 97.10.0 95.80.2 Screw 89.60.0 97.20.0 83.30.7 89.60.5 94.50.4 95.60.4 89.80.8 90.00.7 95.20.3 95.60.6 90.80.2 91.31.0 96.00.5 96.30.5 Tile 77.60.0 96.20.0 84.11.1 94.10.5 96.30.2 97.70.1 87.70.2 94.40.2 96.50.1 97.80.1 88.90.3 94.60.1 96.60.1 97.80.1 Toothbrush 86.90.0 93.20.0 97.30.3 97.30.4 97.80.1 99.10.1 97.70.3 97.50.2 98.10.1 99.20.1 98.40.2 98.40.4 98.40.5 99.20.1 Transistor 74.70.0 69.20.0 90.22.8 84.92.7 85.01.8 87.32.6 92.32.1 89.60.9 88.31.0 88.60.8 94.02.7 90.71.4 88.51.2 90.30.6 Wood 93.40.0 96.00.0 90.70.4 92.70.9 94.61.0 95.71.2 91.90.1 93.20.7 95.30.4 96.40.3 92.20.1 93.50.3 95.40.2 96.40.4 Zipper 91.60.0 98.20.0 93.90.8 97.40.4 93.90.8 98.40.3 95.40.3 98.00.1 94.10.7 98.30.1 96.10.2 98.10.1 94.20.4 98.50.1 Mean 85.10.0 90.90.0 89.30.9 92.01.0 95.20.5 95.80.1 91.30.7 93.30.6 96.00.3 96.30.1 92.60.7 94.30.5 96.20.3 96.80.1
[…]

### v3

@@ A.4 Further results in the one-for-all paradigm
| CLIPFUSION | 96.4 | 98.2 | 88.3 | 90.0 | |
A.5 Limitations of few-shot training
Training-based methods inherently suffer from epoch selection sensitivity. Figure 4(a) shows that the epoch at which PromptAD performs best differs significantly across object categories. In conventional training setups, such epoch selection is guided by a validation set. However, in few-shot anomaly detection, where only one to four images are available, it is practically hard to reserve any for a separate validation set. The authors of PromptAD circumvent this by using the test set for validation in their official implementation.555https://github.com/FuNz-0/PromptAD However, test sets are not accessible during model training in real-world applications. In contrast, CLIPFUSION requires no additional training and uses few-shot examples solely for feature extraction, making validation sets and epoch selection unnecessary.
Figure 4(b) reveals that even when the training epoch is fixed, performance varies considerably across different random seeds. This highlights the instability and seed sensitivity of few-shot training, making it difficult to ensure consistent and reproducible results without a validation set.
A.6 Effect of weights between the CLIP and diffusion models
We analyze the effect of varying the weights assigned to the CLIP and diffusion models. Recall that and denote the weights of the CLIP and the diffusion model, respectively. By changing , we adjust their relative contributions and evaluate the performance. As shown in Table 15, segmentation achieves the best performance among the tested alpha values at , while classification performed best at .
| Task | Weight | Dataset | ||
| CLIP | Diff | MVTec-AD | VisA | |
| Seg. | 1 | 0 | 94.7 | 93.7 |
| 0.75 | 0.25 | 95.4 | 96.1 | |
| 0.5 | 0.5 | 95.7 | 97.2 | |
| 0.25 | 0.75 | 95.8 | 97.3 | |
| 0 | 1 | 95.1 | 96.5 | |
| Cls. | 1 | 0 | 93.5 | 82.2 |
| 0.75 | 0.25 | 95.4 | 86.5 | |
| 0.5 | 0.5 | 95.0 | 86.4 | |
| 0.25 | 0.75 | 92.7 | 83.9 | |
| 0 | 1 | 87.2 | 78.9 | |
A.7 Failure cases
Figure 5 provides representative failure cases. In Figure 5(a), although the anomaly lies in the outer sheath of the cables, the diffusion-based model over-activates on the dense and repetitive wire patterns, failing to highlight the true defect. Figure 5(b) shows a logical anomaly, where the absence of a transistor itself constitutes the defect. Lacking this contextual understanding, the models struggle to determine where to attend, resulting in mislocalization.
A.8 Results on different backbones
Table 16 reports the performance of our framework with different diffusion backbones. We observe that the results are consistent across backbones. This indicates that our method is robust to the choice of diffusion backbone.
| Backbone | Task | Dataset | |
| MVTec-AD | VisA | ||
| Stable Diffusion v1.5 | Seg. | 95.8 | 96.7 |
| Cls. | 93.5 | 86.4 | |
| Stable Diffusion v2.0 | Seg. | 95.8 | 97.3 |
| Cls. | 95.4 | 86.5 | |
| Stable Diffusion v2.1 | Seg. | 95.8 | 97.3 |
| Cls. | 94.6 | 86.5 | |
A.9 Pseudocode
A.10 Detailed quantitative results
| Category | k=0 | k=1 | k=2 | k=4 | ||||||||||
| WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | |
| Bottle | 99.20.0 | 100.00.0 | 97.40.7 | 99.40.4 | 98.20.9 | 100.00.0 | 98.51.0 | 99.20.3 | 99.30.3 | 100.00.0 | 98.80.2 | 99.20.3 | 99.30.4 | 100.00.0 |
| Cable | 86.50.0 | 87.30.0 | 57.74.6 | 88.84.2 | 88.91.9 | 95.00.9 | 62.35.9 | 91.02.7 | 88.40.7 | 96.30.4 | 70.06.1 | 91.02.7 | 90.90.9 | 96.70.4 |
| Capsule | 72.90.0 | 88.10.0 | 57.77.3 | 67.82.9 | 72.36.8 | 83.66.9 | 64.33.0 | 72.87.0 | 77.38.8 | 89.02.7 | 65.22.5 | 72.87.0 | 82.38.9 | 91.21.6 |
| Carpet | 100.00.0 | 99.90.0 | 96.61.0 | 95.30.8 | 99.80.3 | 100.00.0 | 97.80.5 | 96.60.5 | 99.80.3 | 100.00.1 | 97.90.4 | 96.60.5 | 100.00.0 | 100.00.0 |
| Grid | 98.80.0 | 99.90.0 | 54.26.7 | 63.610.3 | 99.50.3 | 99.40.4 | 67.24.2 | 67.78.3 | 99.40.2 | 99.60.4 | 68.13.8 | 67.78.3 | 99.60.1 | 99.50.4 |
| Hazelnut | 93.90.0 | 96.90.0 | 88.32.6 | 88.32.7 | 97.51.4 | 99.30.4 | 90.80.8 | 93.23.8 | 98.30.7 | 99.30.3 | 91.91.2 | 93.23.8 | 98.40.4 | 99.30.5 |
| Leather | 100.00.0 | 99.90.0 | 97.50.7 | 97.30.7 | 99.90.0 | 99.90.0 | 97.50.9 | 97.90.7 | 99.90.0 | 100.00.0 | 98.50.2 | 97.90.7 | 100.00.0 | 100.00.0 |
| Metal nut | 97.10.0 | 98.30.0 | 53.03.8 | 73.42.9 | 98.70.8 | 99.80.3 | 54.83.8 | 77.78.5 | 99.40.2 | 99.90.2 | 60.75.2 | 77.78.5 | 99.50.2 | 100.00.0 |
| Pill | 79.10.0 | 77.60.0 | 61.33.8 | 81.92.8 | 91.22.1 | 96.10.4 | 59.16.4 | 82.92.9 | 92.30.7 | 96.20.2 | 54.92.7 | 82.92.9 | 92.81.0 | 96.40.4 |
| Screw | 83.30.0 | 77.30.0 | 55.02.5 | 44.44.6 | 86.40.9 | 72.58.5 | 54.04.4 | 49.03.8 | 86.02.1 | 73.55.1 | 50.04.1 | 49.03.8 | 87.91.2 | 82.04.5 |
| Tile | 100.00.0 | 99.10.0 | 92.22.2 | 99.00.9 | 99.90.0 | 99.90.2 | 93.31.1 | 98.51.0 | 99.90.2 | 99.90.1 | 93.10.6 | 98.51.0 | 99.90.1 | 100.00.0 |
| Toothbrush | 87.50.0 | 96.90.0 | 82.51.2 | 83.33.8 | 92.24.9 | 98.91.6 | 87.64.2 | 85.93.5 | 97.51.6 | 99.10.8 | 89.22.5 | 85.93.5 | 96.72.6 | 98.50.8 |
| Transistor | 88.00.0 | 86.70.0 | 73.36.0 | 78.16.9 | 83.43.8 | 90.91.4 | 72.86.3 | 90.04.3 | 85.31.7 | 91.10.8 | 82.46.5 | 90.04.3 | 85.72.5 | 92.91.0 |
| Wood | 99.40.0 | 96.70.0 | 96.11.2 | 97.80.3 | 99.90.1 | 98.90.3 | 96.90.5 | 98.30.6 | 99.90.1 | 98.60.2 | 97.00.2 | 98.30.6 | 99.80.3 | 98.80.1 |
| Zipper | 91.50.0 | 94.00.0 | 85.82.7 | 92.30.5 | 88.85.9 | 96.50.7 | 86.32.6 | 94.02.1 | 94.01.4 | 95.32.2 | 88.32.0 | 94.02.1 | 94.50.5 | 96.20.6 |
| Mean | 91.80.0 | 93.20.0 | 76.63.1 | 83.43.0 | 93.12.0 | 95.40.8 | 78.93.1 | 86.33.3 | 94.41.3 | 95.90.3 | 80.42.5 | 88.82.6 | 95.21.3 | 96.80.2 |
| Category | k=0 | k=1 | k=2 | k=4 | ||||||||||
| WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | |
| Bottle | 99.80.0 | 100.00.0 | 99.20.2 | 99.80.1 | 99.40.3 | 100.00.0 | 99.60.3 | 99.80.1 | 99.80.1 | 100.00.0 | 99.70.0 | 99.80.1 | 99.80.1 | 100.00.0 |
| Cable | 91.20.0 | 92.70.0 | 64.93.8 | 93.82.2 | 93.21.1 | 97.10.5 | 69.66.6 | 95.11.3 | 92.90.6 | 97.80.3 | 76.15.6 | 97.10.7 | 94.40.3 | 98.10.2 |
| Capsule | 91.50.0 | 97.20.0 | 86.92.2 | 89.42.0 | 91.62.7 | 95.92.2 | 88.40.8 | 91.02.9 | 93.33.6 | 97.40.8 | 87.80.8 | 94.91.1 | 95.13.3 | 98.00.4 |
| Carpet | 100.00.0 | 100.00.0 | 99.00.2 | 98.70.2 | 99.90.1 | 100.00.0 | 99.40.1 | 99.00.1 | 99.90.1 | 100.00.0 | 99.40.1 | 98.80.2 | 100.00.0 | 100.00.0 |
| Grid | 99.60.0 | 100.00.0 | 75.03.3 | 81.14.9 | 99.90.1 | 99.80.1 | 82.52.3 | 84.14.0 | 99.80.1 | 99.90.1 | 83.01.8 | 86.44.0 | 99.90.0 | 99.80.1 |
| Hazelnut | 96.90.0 | 98.30.0 | 93.31.7 | 92.92.2 | 98.60.7 | 99.60.2 | 94.10.5 | 96.02.0 | 99.10.4 | 99.60.1 | 94.80.6 | 97.01.2 | 99.10.2 | 99.60.3 |
| Leather | 100.00.0 | 100.00.0 | 99.20.2 | 99.10.2 | 100.00.0 | 100.00.0 | 99.20.3 | 99.30.2 | 100.00.0 | 100.00.0 | 99.60.1 | 99.60.1 | 100.00.0 | 100.00.0 |
| Metal nut | 99.30.0 | 99.60.0 | 82.02.7 | 91.01.1 | 99.70.2 | 99.90.1 | 82.21.4 | 92.34.0 | 99.90.0 | 100.00.0 | 85.51.7 | 97.02.6 | 99.90.1 | 100.00.0 |
| Pill | 95.70.0 | 94.80.0 | 88.31.3 | 96.50.6 | 98.30.5 | 99.30.1 | 87.92.6 | 96.60.7 | 98.60.1 | 99.30.0 | 87.01.2 | 96.90.4 | 98.60.2 | 99.40.1 |
| Screw | 93.10.0 | 89.30.0 | 78.11.0 | 71.42.3 | 94.20.6 | 87.54.7 | 77.31.3 | 72.93.4 | 94.11.5 | 87.43.2 | 75.72.8 | 71.81.9 | 94.90.8 | 93.02.3 |
| Tile | 100.00.0 | 99.60.0 | 97.20.7 | 99.60.3 | 100.00.0 | 100.00.1 | 97.60.4 | 99.40.4 | 100.00.1 | 100.00.0 | 97.60.2 | 99.60.1 | 100.00.0 | 100.00.0 |
| Toothbrush | 95.60.0 | 98.80.0 | 93.70.5 | 93.51.4 | 96.72.0 | 99.60.5 | 95.21.6 | 94.11.4 | 99.00.6 | 99.70.3 | 95.80.7 | 94.80.7 | 98.71.1 | 99.50.2 |
| Transistor | 87.10.0 | 84.70.0 | 66.27.5 | 77.75.5 | 79.04.0 | 89.81.5 | 69.06.5 | 89.33.9 | 80.72.3 | 89.90.7 | 77.68.4 | 84.59.0 | 80.73.2 | 92.31.3 |
| Wood | 99.80.0 | 99.00.0 | 98.80.3 | 99.30.1 | 100.00.0 | 99.70.1 | 99.00.1 | 99.50.2 | 100.00.0 | 99.60.1 | 99.10.0 | 99.50.2 | 99.90.1 | 99.70.0 |
| Zipper | 97.50.0 | 97.90.0 | 95.50.9 | 97.20.3 | 96.81.8 | 98.90.3 | 95.41.0 | 97.81.0 | 98.30.4 | 98.31.0 | 96.20.8 | 99.10.7 | 98.50.2 | 98.70.2 |
| Mean | 96.50.0 | 96.80.0 | 88.11.7 | 92.21.5 | 96.50.9 | 97.80.4 | 89.31.7 | 93.81.7 | 97.00.7 | 97.90.2 | 90.51.6 | 94.51.5 | 97.30.6 | 98.50.1 |
| Category | k=0 | k=1 | k=2 | k=4 | ||||||||||
| WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | |
| Bottle | 89.50.0 | 94.10.0 | 96.10.5 | 97.90.1 | 97.50.2 | 97.30.2 | 96.90.1 | 98.10.0 | 97.70.1 | 97.50.0 | 97.10.1 | 98.20.0 | 97.80.0 | 97.70.1 |
| Cable | 77.00.0 | 73.20.0 | 88.41.2 | 95.50.8 | 93.80.6 | 91.30.2 | 90.00.8 | 96.40.3 | 94.30.4 | 92.20.3 | 92.10.4 | 97.50.3 | 94.90.1 | 92.90.4 |
| Capsule | 86.90.0 | 94.60.0 | 94.50.6 | 95.60.4 | 94.60.8 | 97.90.4 | 95.20.5 | 96.50.4 | 96.40.3 | 98.40.2 | 96.20.4 | 96.80.6 | 96.20.5 | 98.50.1 |
| Carpet | 95.40.0 | 99.30.0 | 97.80.2 | 98.40.1 | 99.40.0 | 99.40.1 | 98.20.0 | 98.50.1 | 99.30.0 | 99.50.0 | 98.40.0 | 98.60.1 | 99.30.0 | 99.50.0 |
| Grid | 82.20.0 | 98.90.0 | 70.22.8 | 58.84.9 | 96.81.0 | 98.80.2 | 70.82.0 | 62.63.2 | 97.70.8 | 98.90.4 | 77.01.8 | 69.41.3 | 98.00.2 | 99.20.1 |
| Hazelnut | 94.30.0 | 96.70.0 | 95.40.7 | 95.80.6 | 98.50.2 | 98.20.1 | 96.80.3 | 96.30.6 | 98.70.1 | 98.20.1 | 97.20.2 | 97.60.1 | 98.80.0 | 98.70.1 |
| Leather | 96.70.0 | 99.20.0 | 98.50.1 | 98.80.2 | 99.30.0 | 99.30.0 | 98.70.1 | 99.00.1 | 99.30.0 | 99.30.0 | 98.80.0 | 99.10.0 | 99.30.0 | 99.40.0 |
| Metal nut | 61.00.0 | 74.20.0 | 74.61.1 | 89.31.4 | 90.00.6 | 86.21.2 | 80.32.1 | 94.61.4 | 91.40.4 | 89.70.9 | 82.73.9 | 95.91.8 | 92.90.4 | 91.30.8 |
| Pill | 80.00.0 | 83.00.0 | 84.81.0 | 93.11.1 | 96.40.3 | 95.50.2 | 87.30.7 | 94.20.3 | 97.00.2 | 95.60.2 | 88.90.5 | 94.80.4 | 97.10.0 | 95.80.2 |
| Screw | 89.60.0 | 97.20.0 | 83.30.7 | 89.60.5 | 94.50.4 | 95.60.4 | 89.80.8 | 90.00.7 | 95.20.3 | 95.60.6 | 90.80.2 | 91.31.0 | 96.00.5 | 96.30.5 |
| Tile | 77.60.0 | 96.20.0 | 84.11.1 | 94.10.5 | 96.30.2 | 97.70.1 | 87.70.2 | 94.40.2 | 96.50.1 | 97.80.1 | 88.90.3 | 94.60.1 | 96.60.1 | 97.80.1 |
| Toothbrush | 86.90.0 | 93.20.0 | 97.30.3 | 97.30.4 | 97.80.1 | 99.10.1 | 97.70.3 | 97.50.2 | 98.10.1 | 99.20.1 | 98.40.2 | 98.40.4 | 98.40.5 | 99.20.1 |
| Transistor | 74.70.0 | 69.20.0 | 90.22.8 | 84.92.7 | 85.01.8 | 87.32.6 | 92.32.1 | 89.60.9 | 88.31.0 | 88.60.8 | 94.02.7 | 90.71.4 | 88.51.2 | 90.30.6 |
| Wood | 93.40.0 | 96.00.0 | 90.70.4 | 92.70.9 | 94.61.0 | 95.71.2 | 91.90.1 | 93.20.7 | 95.30.4 | 96.40.3 | 92.20.1 | 93.50.3 | 95.40.2 | 96.40.4 |
| Zipper | 91.60.0 | 98.20.0 | 93.90.8 | 97.40.4 | 93.90.8 | 98.40.3 | 95.40.3 | 98.00.1 | 94.10.7 | 98.30.1 | 96.10.2 | 98.10.1 | 94.20.4 | 98.50.1 |
| Mean | 85.10.0 | 90.90.0 | 89.30.9 | 92.01.0 | 95.20.5 | 95.80.1 | 91.30.7 | 93.30.6 | 96.00.3 | 96.30.1 | 92.60.7 | 94.30.5 | 96.20.3 | 96.80.1 |
| Category | k=0 | k=1 | k=2 | k=4 | ||||||||||
| WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | PaDiM | PatchCore | WinCLIP | CLIPFUSION | |
| Bottle | 76.40.0 | 88.40.0 | 89.80.8 | 93.50.3 | 91.20.4 | 93.70.4 | 91.70.2 | 93.90.3 | 91.80.3 | 94.10.1 | 92.20.2 | 94.00.2 | 91.60.2 | 94.50.3 |
| Cable | 42.90.0 | 71.50.0 | 59.13.2 | 84.71.0 | 72.52.3 | 79.70.8 | 66.52.8 | 88.50.9 | 74.72.3 | 82.01.0 | 74.21.8 | 91.70.6 | 77.01.1 | 84.41.0 |
| Capsule | 62.10.0 | 88.50.0 | 80.02.0 | 83.90.9 | 85.62.7 | 92.51.1 | 82.32.1 | 86.61.0 | 90.60.6 | 94.10.2 | 85.71.3 | 87.81.9 | 90.11.5 | 94.70.5 |
| Carpet | 84.10.0 | 98.20.0 | 92.90.3 | 93.30.3 | 97.40.4 | 98.30.1 | 93.90.2 | 93.70.4 | 97.30.3 | 98.40.1 | 94.40.2 | 93.90.4 | 97.00.2 | 98.50.1 |
| Grid | 57.00.0 | 95.70.0 | 41.24.6 | 21.79.5 | 90.52.7 | 96.11.1 | 45.13.6 | 23.73.8 | 92.82.5 | 97.10.9 | 55.53.4 | 30.44.6 | 93.60.6 | 96.90.7 |
| Hazelnut | 81.60.0 | 93.80.0 | 85.71.9 | 88.31.3 | 93.70.9 | 94.30.3 | 89.40.9 | 89.81.3 | 94.20.3 | 94.70.5 | 90.40.7 | 92.00.3 | 94.20.3 | 95.60.6 |
| Leather | 91.10.0 | 97.70.0 | 95.60.2 | 95.21.0 | 98.60.0 | 97.80.1 | 96.20.2 | 95.90.3 | 98.30.4 | 97.70.1 | 96.30.1 | 96.40.1 | 98.00.4 | 97.60.1 |
| Metal nut | 31.80.0 | 74.10.0 | 38.11.6 | 66.72.9 | 84.71.1 | 80.11.6 | 48.25.0 | 79.64.2 | 86.70.8 | 85.81.6 | 54.08.8 | 83.85.5 | 89.40.1 | 87.61.1 |
| Pill | 65.00.0 | 87.60.0 | 78.90.6 | 89.51.6 | 93.50.2 | 93.90.3 | 84.30.4 | 91.60.5 | 94.50.2 | 94.20.2 | 86.60.4 | 92.50.4 | 94.60.3 | 94.50.1 |
| Screw | 68.50.0 | 89.50.0 | 51.61.7 | 68.11.3 | 82.31.1 | 84.00.8 | 69.52.1 | 69.02.1 | 84.10.5 | 83.61.5 | 72.30.8 | 72.43.1 | 86.31.8 | 85.91.1 |
[…]